# On the convergence of bootstrap current to the Shaing-Callen limit in stellarators

### v1

@@ 4.1 Long field lines
The resulting normalized bootstrap coefficients (43) are shown in Fig. 18. They are weakly dependent on $\iota$ for collisionalities $\nu_{\ast}>10^{-5}$ (in the first hill), and follow this trend down to $\nu_{\ast}=10^{-6}$ in the second hill, with the only curve, $\iota=0.43$, departing from the other four. As one can see from the $\iota$ scan in this figure, this transition is smooth, in contrast to the Shaing-Callen limit which show well known “bootstrap resonances” briefly discussed in Section 4.4.
@@
Since the distribution function in the trapped particle domain is almost independent of pitch parameter (equivalently of $\eta$), the plotted quantity characterizes well the radial particle flux. Obviously, this function is stellarator anti-symmetric, $g_{(3)}(\vartheta,\varphi,0)=-g_{(3)}(\vartheta,\varphi,0)$, and retains all other relevant features seen for the short field line examples in Section 2.3. Namely, “hot spots” around the global maximum ($\vartheta=\pi$, $\varphi=\pi/3$) where $g_{(3)}$ is the largest correspond to the contact region with the boundary layer (see the left Fig. 7). These spots are responsible for the tokamak-like off-set studied by Helander et al. (2011). One can clearly see in Fig. 19 also the off-set domains (blue and yellow) in the ripples adjacent to the global maximum (mainly contributing to the off-set at this collisionality) as well as widely extending “main region” (green). Thus, this case is rather similar to Fig. 6 for $\nu_{\ast}=10^{-4}$. Remarkably, results for two different $\iota$ are almost indistinguishable, which explains close values of $\lambda_{bB}$ in Fig. 18 at this collisionality. Since the radial guiding center velocity $v_{g}^{r}$ is stellarator antisymmetric as well, the resulting normalized radial trapped particle flux distribution over the angles $\gamma$ given with good accuracy by $\gamma(\vartheta,\varphi)={\cal V}(\vartheta,\varphi)g_{(3)}(\vartheta,\varphi,0)$ where ${\cal V}$ is the sub-integrand in Eq. (114), is stellarator-symmetric. It is shown for $\nu_{\ast}=3\cdot 10^{-4}$ in Fig. 20 where the contribution of the off-set domain ($42<\vartheta<$ 4) is positive what agrees with the first hill in Fig. 18. Note that direct contribution of the boundary layer is already small at this collisionality which is also seen from vanishing hot spots in Fig. 20.
@@ 4.2 Asymptotic behavior at irrational surfaces
For estimation of $\lambda_{bB}$ trend with $\nu_{\ast}\rightarrow 0$ we analyze a nearly periodic pattern seen in right plots of Figs. 24 and 25 (such pattern is seen also in the left plot of Fig. 23 while the right plot shows the overlay of two patterns). This pattern is formed if the field line starting in close vicinity of global maximum point $(\vartheta_{\rm max},\varphi_{\rm max})$, i.e. from the point $(\vartheta_{0},\varphi_{0})=(\vartheta_{\rm max}+\Delta\vartheta_{0},\varphi_{\rm max})$ where $\Delta\vartheta_{0}\ll\delta\vartheta$ and $\delta\vartheta$ is the poloidal boundary layer width (width of “hot spots” in Figs. 19 and 21), returns after $N$ turns back to this vicinity, $(\vartheta_{N},\varphi_{N})=(\vartheta_{\rm max}+\Delta\vartheta_{N},\varphi_{\rm max})$ with $\Delta\vartheta_{N}\ll\delta\vartheta$. The off-set pattern is formed if within these $N$ turns at least one approximate match occurs at some $(\vartheta,\varphi)=(\vartheta_{\rm max}+\Delta\vartheta,\varphi_{\rm max})$ with $\Delta\vartheta_{0}\ll\Delta\vartheta\leq\delta\vartheta$. By the symmetry argument (one can make $N$ turns backwards), such matches come in pairs thus forming one “main region” and at least two off-set domains with positive and negative off-set respectively (see the left plot in Fig. 23 and the right plots in Figs. 24 and 25 where six off-set domains are formed for $\iota=0.435$).
@@
where second derivatives correspond to the global maximum point, and mixed derivative there is zero for our example (59), we can relate the distance $\Delta\eta$ between class transition boundary and trapped-passing boundary $\eta=\eta_{b}=1/B_{\rm max}$ to the poloidal distance $\Delta\vartheta$ between the respective local maximum and the global maximum as
@@
The term $\iota^{2}\beta\sim 1$ results here from the small tilt from $\varphi=\text{const}$ of the ${\bf h}\cdot\nabla B=0$ line seen in Fig. 19 as a boundary between positive and negative off-set domains (and the main region). Assuming that rational numbers $10<m/n<$ 1 with $n<N_{\max}$ are distributed uniformly in this interval (which is generally not true but gives a correct order of magnitude estimate) we obtain for the best match at some $n=N$ that
@@
Thus, field line starting at global maximum will be displaced from this maximum after $N$ toroidal turns by $\Delta\vartheta_{N}\sim 2\pi/N$ and, according to (117), pass through the local maximum with respective class boundary $\eta_{c}$ displacement $\Delta\eta_{N}=\eta_{c}-\eta_{b}$ given by
@@
where $B_{0}=B_{\rm max}$, and constant (76) has been estimated as $C_{0}\sim RB_{0}^{2}/(\nu_{\ast}\varepsilon_{t}^{1/2})$. Note that up to a factor 2 scaling (121) is reproduced by the relation between color ranges in Figs. 19 and 21.
@@
with $\eta=\eta_{\text{loc}}^{(j)}$, this series has no $m=0$ terms due to the last (115) where derivative over $\vartheta_{0}$ is the same with derivative over $\vartheta$. Here, we placed the origin of angles at the global $B$ maximum to avoid cosine harmonics in stellarator symmetric fields. Introducing the number of toroidal turns per off-set domain, $N_{o}\lesssim N$, we present $\iota=\iota_{o}+\delta\iota$ where $\iota_{o}=M_{o}/N_{o}$, $M_{o}$ is number of poloidal turns and $\delta\iota\ll 1$. Field line segment with the off-set domain starts at $(\vartheta,\varphi)=(\Delta\vartheta_{0},0)$ and ends at $(\vartheta,\varphi)=(2\pi M_{o}+\Delta\vartheta_{N},2\pi N_{0})$ where $\Delta\vartheta_{N}=\Delta\vartheta_{0}+2\pi N_{o}\delta\iota$ and banana tips are located at $(\vartheta_{j},\varphi_{j})=(\Delta\vartheta_{0}+\iota\delta\varphi_{0},\delta\varphi_{0})$ and $(\vartheta_{j+1},\varphi_{j+1})=(2\pi M_{o}+\Delta\vartheta_{N}+\iota\delta\varphi_{N},2\pi N_{o}+\delta\varphi_{N})$, where $\delta\varphi_{0,N}\sim\max(|\Delta\vartheta_{0}|,|\Delta\vartheta_{N}||)\ll 1$. Expanding the integrand in (114) around the helical line $\vartheta=\iota_{o}\varphi$ we get
$I_{{\cal V}}^{(j)}$ $=$ $\int\limits_{\varphi_{j}}^{\varphi_{j+1}}{\rm d}\varphi{\cal V}\left(\Delta\vartheta_{0}+\iota\varphi,\varphi\right)$
@@
where ${\cal V}_{b}$ is given by Eq. (115) with $\eta_{\text{loc}}^{(j)}\rightarrow\eta_{b}$, and ${\cal V}^{\prime}$ and ${\cal V}^{\prime\prime}$ denote first and second derivatives of ${\cal V}(\vartheta,\varphi)$ over $\vartheta$. Approximations in Eq. (4.2) ignoring $\delta\varphi_{0,j}$ and $\eta_{\text{loc}}^{(j)}-\eta_{b}$ skip the terms of the order of $\Delta\vartheta_{0,N}^{3}$. In the case of stellarator symmetry, quantities ${\cal V}_{b}$ and ${\cal V}_{b}^{\prime\prime}$ are odd functions of $\varphi-\pi N_{o}$ while ${\cal V}_{b}^{\prime}$ is an even function, what results in
@@
where $\delta_{\text{res}}=1$ if $mM_{o}-nN_{o}=0$ and is zero otherwise, and $\Delta\vartheta_{\text{mid}}=(\Delta\vartheta_{0}+\Delta\vartheta_{N})/2$. Since only the resonant harmonic with $(m,n)=\left(N_{o}k,-M_{o}k\right)$ and $k=1,2,\dots$ contribute to the first term, one can estimate $\sum\sum m{\cal V}_{mn}^{s}\delta_{\text{res}}\sim N_{o}{\cal V}_{N_{o},-M_{o}}^{s}$. For high harmonic indices, $m,n\rightarrow\infty$, spectrum of ${\cal V}$ is determined by the vicinity of the global maximum,
${\cal V}_{mn}^{s}(\eta_{b})\approx\frac{9\beta^{1/2}\eta_{b}^{7/2}}{\sqrt{8}\pi}\left|\frac{\partial^{2}B}{\partial\vartheta^{2}}\right|^{3/2}\frac{m}{(m^{2}+\beta n^{2})^{5/2}},$ (125)
see also Eq. (116). Therefore, contribution of the first term $N_{o}{\cal V}_{N_{o},-M_{o}}^{s}\sim N_{o}^{-3}$ is negligible small compared to that of the second term which does not scale with $N_{o}$. Thus we obtain the scaling $I_{{\cal V}}^{(j)}\propto(\Delta\vartheta_{N}^{2}-\Delta\vartheta_{0}^{2})\propto N_{o}^{-2}\propto N^{-2}$ and respective estimates
@@
where we assumed the maximum distribution function off-set (121), used Eqs. (120) and assumed $\varepsilon_{M}\sim\varepsilon_{t}$ and $\beta\sim 1$. Thus, off-set of bootstrap coefficient $\lambda_{\text{off}}$ in stellarator-symmetric fields does not decrease with reducing collisionality but has an aperiodic oscillatory behaviour with oscillation amplitude $\lambda_{\text{off}}^{\text{max}}$ of the order of $\lambda_{bB}$ for the equivalent tokamak (Boozer and Gardner, 1990; Beidler et al., 2011),
@@
see Fig. 18 where the factor $1.46\iota^{-1}\approx 3.4$ agrees well with the relative off-set amplitude. The off-set structure corresponding to $N$ toroidal turns is replaced by the structure with larger number of turns if collisionality is reduced by a significant factor such that condition $\Delta\eta_{N}\lesssim\delta\eta$ changes to $\Delta\eta_{N}\gg\delta\eta$ where $\Delta\eta_{N}$ and $\delta\eta$ are given by Eqs. (118) and (119), respectively. In this case, the previous off-set vanishes exponentially and is replaced by the next one. This means that the overall off-set changes if $\log\nu_{\ast}$ changes by a value of the order one, which is roughly the trend in Fig. 18.

### v3

@@ 4.1 Long field lines
The resulting normalized bootstrap coefficients (43) are shown in Fig. 20. They are weakly dependent on $\iota$ for collisionalities $\nu_{\ast}>10^{-5}$ (in the first hill), and follow this trend down to $\nu_{\ast}=10^{-6}$ in the second hill, with the only curve, $\iota=0.43$, departing from the other four. As one can see from the $\iota$ scan in this figure, this transition is smooth, in contrast to the Shaing-Callen limit which shows well known “bootstrap resonances” briefly discussed in Section 4.4.
@@
Since the distribution function in the trapped particle domain is almost independent of pitch parameter (equivalently of $\eta$), the plotted quantity characterizes well the radial particle flux. Obviously, this function is stellarator anti-symmetric, $g_{(3)}(\vartheta,\varphi,0)=-g_{(3)}(-\vartheta,-\varphi,0)$, and retains all other relevant features seen for the short field line examples in Section 2.3. Namely, “hot spots” around the global maximum ($\vartheta=\pi$, $\varphi=\pi/3$) where $g_{(3)}$ is the largest correspond to the contact region with the boundary layer (see the left Fig. 7). These spots are responsible for the tokamak-like off-set studied by Helander et al. (2011). One can clearly see in Fig. 21 also the off-set domains (blue and yellow) in the ripples adjacent to the global maximum (mainly contributing to the off-set at this collisionality) as well as widely extending “main region” (green). Thus, this case is rather similar to Fig. 6 for $\nu_{\ast}=10^{-4}$. Remarkably, results for two different $\iota$ are almost indistinguishable, which explains close values of $\lambda_{bB}$ in Fig. 20 at this collisionality. Since the radial guiding center velocity $v_{g}^{r}$ is stellarator antisymmetric as well, the resulting normalized radial trapped particle flux distribution over the angles $\gamma$ given with good accuracy by $\gamma(\vartheta,\varphi)={\cal V}(\vartheta,\varphi)g_{(3)}(\vartheta,\varphi,0)$ where ${\cal V}$ is the sub-integrand in Eq. (111), is stellarator-symmetric. It is shown for $\nu_{\ast}=3\cdot 10^{-4}$ in Fig. 22 where the contribution of the off-set domain ($42<\vartheta<$ 4) is positive what agrees with the first hill in Fig. 20. Note that direct contribution of the boundary layer is already small at this collisionality which is also seen from vanishing hot spots in Fig. 22.
@@ 4.2 Asymptotic behavior at irrational surfaces
For estimation of $\lambda_{bB}$ trend with $\nu_{\ast}\rightarrow 0$ we analyze a nearly periodic pattern seen in right plots of Figs. 26 and 27 (such pattern is seen also in the left plot of Fig. 25 while the right plot shows the overlay of two patterns). This pattern is formed if the field line starting in close vicinity of global maximum point $(\vartheta_{\rm max},\varphi_{\rm max})$, i.e. from the point $(\vartheta_{0},\varphi_{0})=(\vartheta_{\rm max}+\Delta\vartheta_{0},\varphi_{\rm max})$ where $\Delta\vartheta_{0}\ll\delta\vartheta$ and $\delta\vartheta$ is the poloidal boundary layer width (width of “hot spots” in Figs. 21 and 23), returns after $N$ turns back to this vicinity, $(\vartheta_{N},\varphi_{N})=(\vartheta_{\rm max}+\Delta\vartheta_{N},\varphi_{\rm max})$ with $\Delta\vartheta_{N}\ll\delta\vartheta$. The off-set pattern is formed if within these $N$ turns at least one approximate match occurs at some $(\vartheta,\varphi)=(\vartheta_{\rm max}+\Delta\vartheta,\varphi_{\rm max})$ with $\Delta\vartheta_{0}\ll\Delta\vartheta\leq\delta\vartheta$. By the symmetry argument (one can make $N$ turns backwards), such matches come in pairs thus forming one “main region” and at least two off-set domains with positive and negative off-set respectively (see the left plot in Fig. 25 and the right plots in Figs. 26 and 27 where six off-set domains are formed for $\iota=0.435$).
@@
where second derivatives correspond to the global maximum point, and mixed derivative there is zero for our example (55), we can relate the distance $\Delta\eta$ between class transition boundary $\eta=\eta_{c}=1/B^{\text{loc}}_{\text{max}}$ and trapped-passing boundary $\eta=\eta_{b}=1/B_{\rm max}$ to the poloidal distance $\Delta\vartheta$ between the respective local maximum and the global maximum as
@@
We have expressed here the toroidal distance between these maxima as $\Delta\varphi=-\iota\beta\Delta\vartheta$ which follows from the local maximum condition ${\bf h}\cdot\nabla B=0$ and Eq. (113). Assuming that rational numbers $10<m/n<$ 1 with $n<N_{\max}$ are distributed uniformly in this interval (which is generally not true but gives a correct order of magnitude estimate) we obtain for the best match at some $n=N$ that
@@
Thus, a field line starting at a global maximum will be displaced from this maximum after $N$ toroidal turns by $\Delta\vartheta_{N}\sim 2\pi/N$ and, according to (114), pass through the local maximum with respective class boundary $\eta_{c}$ displacement $\Delta\eta_{N}=\eta_{c}-\eta_{b}$ given by
@@
where $B_{0}=B_{\rm max}$, and constant (72) has been estimated as $C_{0}\sim RB_{0}^{2}/(\nu_{\ast}\varepsilon_{t}^{1/2})$. Note that up to a factor 2 scaling (118) is reproduced by the relation between color ranges in Figs. 21 and 23 and is in good agreement with collisionality dependence of the span $g_{\text{off}}^{\text{(sp)}}$ of the generalized Spitzer function which corresponds to the maximum off-set at $\varphi=0$ line (in particular, in Figs. 21 and 23), see Fig. 28.
Figure 28: Span of the generalized Spitzer function $g_{\text{off}}^{\text{(sp)}}\equiv\left(\max\left(g_{\text{(3)}}\right)-\min\left(g_{\text{(3)}}\right)\right)/2$ for standing particles, $\lambda=0$, at $\varphi=0$ as function of collisionality (case $\iota=0.435$ in Fig. 20). Scalings $\nu_{\ast}^{-3/5}$, $1/\nu_{\ast}$ and $1/\sqrt{\nu_{\ast}}$ are shown by black dashed, red and green dotted lines, respectively.
@@
with $\eta=\eta_{\text{loc}}^{(j)}$, this series has no $m=0$ terms due to the last (112) where the derivative over $\vartheta_{0}$ is the same as the derivative over $\vartheta$. Here, we placed the origin of angles at the global $B$ maximum to avoid cosine harmonics in stellarator symmetric fields. Introducing the number of toroidal turns per off-set domain, $N_{o}\lesssim N$, we present $\iota=\iota_{o}+\delta\iota$ where $\iota_{o}=M_{o}/N_{o}$, $M_{o}$ is the number of poloidal turns and $\delta\iota\ll 1$. Field line segment with the off-set domain starts at $(\vartheta,\varphi)=(\Delta\vartheta_{0},0)$ and ends at $(\vartheta,\varphi)=(2\pi M_{o}+\Delta\vartheta_{N},2\pi N_{0})$ where $\Delta\vartheta_{N}=\Delta\vartheta_{0}+2\pi N_{o}\delta\iota$, and the tips of the transient banana orbit are located at $\left(\vartheta_{(j)}^{-},\varphi_{(j)}^{-}\right)=\left(\Delta\vartheta_{0}+\iota\delta\varphi_{0},\delta\varphi_{0}\right)$ and $\left(\vartheta_{(j)}^{+},\varphi_{(j)}^{+}\right)=\left(2\pi M_{o}+\Delta\vartheta_{N}+\iota\delta\varphi_{N},2\pi N_{o}+\delta\varphi_{N}\right)$, where $\delta\varphi_{0,N}\sim\max(|\Delta\vartheta_{0}|,|\Delta\vartheta_{N}||)\ll 1$ are toroidal shifts of those banana tips from the global maximum points, see Fig. 29. Replacing in $I_{{\cal V}}^{(j)}$, Eq. (111), the integrand ${\cal V}$ with ${\cal V}_{b}$ which is given by the same Eq. (112) with the replacement $\eta_{\text{loc}}^{(j)}\rightarrow\eta_{b}$, extending there the integration interval to the whole off-set segment $0<\varphi<2\pi N_{o}$ and thus ignoring the shifts $\delta\varphi_{0,N}$, and expanding ${\cal V}_{b}$ around the helical line $\vartheta=\iota_{o}\varphi$ we get
$I_{{\cal V}}^{(j)}$ $=$ $\int\limits_{\varphi_{(j)}^{-}}^{\varphi_{(j)}^{+}}{\rm d}\varphi\;{\cal V}\left(\Delta\vartheta_{0}+\iota\varphi,\varphi\right)$
@@
where ${\cal V}_{b}^{\prime}$ and ${\cal V}_{b}^{\prime\prime}$ denote first and second derivatives of ${\cal V}_{b}(\vartheta,\varphi)$ over $\vartheta$. First two approximations in Eq. (4.2) skip negligible small terms of the order of $\Delta\vartheta_{0,N}^{3}$.
Figure 29: Field line segment containing the off-set domain (blue dotted) covered by transient banana orbit (red) in case the lower local field maximum is on the right, $\eta_{\text{loc}}^{(j)}=\eta_{j+1}$. Dashed black ellipses correspond to the contour $\eta_{\text{loc}}^{(j)}B(\vartheta,\varphi)=1$.
To check that we notice that derivative $\partial B/\partial\vartheta_{0}$ in Eq. (112) vanishes at the global maximum and, therefore, scales in the additional integration intervals $0<\varphi<\varphi_{(j)}^{-}$ and $\varphi_{(j)}^{+}<\varphi<2\pi N_{o}$ as $\partial B/\partial\vartheta_{0}\propto\Delta\vartheta_{0,N}$. The same scaling is true there for the square root term $(1-\eta_{b}B)^{1/2}\propto\Delta\vartheta_{0,N}$. Thus, due to the scaling ${\cal V}\propto\Delta\vartheta_{0,N}^{2}$ in those intervals whose size is $|\delta\varphi_{0,N}|$, their overall contribution has a cubic scaling with $\Delta\vartheta_{0,N}$. The difference $\delta{\cal V}={\cal V}-{\cal V}_{b}$ can be estimated as $\delta{\cal V}\propto\eta_{\text{loc}}^{(j)}-\eta_{b}\propto\Delta\vartheta_{0,N}^{2}$ in the whole integration interval including the vicinity of its ends where $\delta{\cal V}\sim{\cal V}_{b}\propto\Delta\vartheta_{0,N}^{2}$. However, in the case of stellarator symmetry, $\delta{\cal V}$ is essentially an odd function of $\varphi-\pi N_{o}$, which does not contribute to the integral, while its even part $\delta{\cal V}^{\text{even}}\propto\Delta\vartheta_{0,N}^{3}$ respectively provides a negligible contribution $\propto\Delta\vartheta_{0,N}^{3}$. Since in the case of stellarator symmetry quantities ${\cal V}_{b}$ and ${\cal V}_{b}^{\prime\prime}$ are odd functions of $\varphi-\pi N_{o}$ while ${\cal V}_{b}^{\prime}$ is an even function, we get
@@
where $\delta_{\text{res}}=1$ if $mM_{o}-nN_{o}=0$ and is zero otherwise, and $\Delta\vartheta_{\text{mid}}=(\Delta\vartheta_{0}+\Delta\vartheta_{N})/2$. Since only the resonant harmonics with $(m,n)=\left(N_{o}k,-M_{o}k\right)$ and $k=1,2,\dots$ contribute to the first term, one can estimate $\sum\sum m{\cal V}_{mn}^{s}\delta_{\text{res}}\sim N_{o}{\cal V}_{N_{o},-M_{o}}^{s}$. For high harmonic indices, $m,n\rightarrow\infty$, spectrum of ${\cal V}$ is determined by the vicinity of the global maximum,
${\cal V}_{mn}^{s}(\eta_{b})\approx\frac{9\beta^{1/2}\eta_{b}^{7/2}}{\sqrt{8}\pi}\left|\frac{\partial^{2}B}{\partial\vartheta^{2}}\right|^{3/2}\frac{mN_{\text{tor}}}{(m^{2}+\beta n^{2})^{5/2}},$ (122)
where $N_{\text{tor}}$ is the number of toroidal field periods and ${\cal V}_{mn}^{s}=0$ for $n$ which are not multiples of $N_{\text{tor}}$, see Appendix A. Therefore, contribution of the first term $N_{o}{\cal V}_{N_{o},-M_{o}}^{s}\sim N_{o}^{-3}$ is negligible small compared to that of the second term which does not scale with $N_{o}$ but contains small $\delta\iota$. Using $2\pi N_{o}\delta\iota=\Delta\vartheta_{N}-\Delta\vartheta_{0}$ and the definition of $\Delta\vartheta_{\text{mid}}$, we obtain the scaling $I_{{\cal V}}^{(j)}\propto(\Delta\vartheta_{N}^{2}-\Delta\vartheta_{0}^{2})\propto N_{o}^{-2}\propto N^{-2}$ and respective estimates
@@
where we assumed the maximum distribution function off-set (118), used Eqs. (117) and assumed $\varepsilon_{M}\sim\varepsilon_{t}$ and $\beta\sim 1$. Thus, off-set of bootstrap coefficient $\lambda_{\text{off}}$ in stellarator-symmetric fields does not decrease with reducing collisionality but has an aperiodic oscillatory behavior with oscillation amplitude $\lambda_{\text{off}}^{\text{max}}$ of the order of $\lambda_{bB}$ for the equivalent tokamak (Boozer and Gardner, 1990; Beidler et al., 2011),
@@
see Fig. 20 where the factor $1.46\iota^{-1}\approx 3.4$ agrees well with the relative off-set amplitude.
The off-set pattern corresponding to $N$ toroidal turns is replaced by the pattern with a larger number of turns if collisionality decreases by a significant factor, and further collisionality decrease triggers another pattern change, etc. Respectively, the off-set of the bootstrap coefficient oscillates aperiodically with changing $\log\nu_{\ast}$ as seen in Fig. 20. This behaviour is due to the structure of the off-set pattern which is formed at the field line segment between two local field maxima $j=0,N$ fulfilling strong alignment conditions $(\Delta\eta_{0},\Delta\eta_{N})\ll\delta\eta$ with $\delta\eta$ given by (116) and which necessarily contains one or more internal local maxima fulfilling a weak alignment condition $\Delta\eta\lesssim\delta\eta$ (they split the segment into off-set domains, e.g., two off-set domains and the main region in the simplest case). With decreasing collisionality, weakly aligned internal maxima become non-aligned, $\Delta\eta\gg\delta\eta$, what makes the variation of the distribution function off-set $g_{\text{off}}$ within the segment and respective contribution of the given pattern to the off-set of $\lambda_{bB}$ exponentially small. Consequently, the whole segment of the old pattern turns into a single off-set domain of the new, longer pattern limited at the ends by even better aligned local maxima such that they stay well aligned when one (or both) maxima at the ends of the former segment become only weakly aligned at some lower collisionality. The above transition is well seen in Figs. 25 and 26 where three complete off-set patterns at higher collisionality (left Fig. 25) turn into three off-set domains of a longer pattern at lower collisionality (left Fig. 26). Note that those domains contribute simultaneously to both off-sets, $g_{\text{off}}^{\text{short}}$ and $g_{\text{off}}^{\text{long}}$ with the first one being dominant at higher and the second one at lower collisionality. A kink on the collisionality dependence of a summary function $g_{\text{off}}^{(sp)}=\max(g_{\text{off}}^{\text{short}},g_{\text{off}}^{\text{long}})$ shown in Fig. 28 is at the transition point $\nu_{\ast}=10^{-4}$ where $g_{\text{off}}^{\text{short}}=g_{\text{off}}^{\text{long}}$. At this point, internal maxima of the long off-set domain are still well aligned, so that $g_{\text{off}}^{\text{long}}$ scaling with $\nu_{\ast}$ is similar to Eq. (105) which includes $1/\nu$ and $1/\sqrt{\nu}$ terms. Note that $g_{\text{off}}^{(sp)}$ in Fig. 28 by definition shows only the largest off-set within the pattern which generally contains more than two off-set domains (see Fig. 26). Therefore, location of $g_{\text{off}}^{(sp)}$ maximum at $\nu_{\ast}$ dependence needs not to coincide with the location of respective $\lambda_{bB}$ maximum, which is at lower $\nu_{\ast}$ in Fig. 20. The reason is that contributions of most off-set domains within the pattern tend to saturate and then vanish with decreasing $\nu_{\ast}$ earlier than $g_{\text{off}}^{(sp)}$ which is determined then by the off-set domain bounded by the best aligned local maxima. One should also note that another kink on $g_{\text{off}}^{(sp)}$ located in Fig. 28 at $\nu_{\ast}=1.7\cdot 10^{-10}$ is not the transition point of different off-set patterns but a point where the off-set for the same pattern changes sign due to the mirroring within trapped-passing boundary layer discussed, in particular, in Section 3.1. As a result, the respective off-set of $\lambda_{bB}$ in Fig. 20 changes sign too (similar to Fig. 2).
According to Eq. (114), better alignment of maxima at the off-set segment ends means smaller poloidal displacements $\Delta\vartheta_{0}$ and $\Delta\vartheta_{N}$ of these ends, which requires a decrease with the pattern length $N$ of the difference $|\Delta\vartheta_{N}-\Delta\vartheta_{0}|=2\pi|\iota N_{p}-M|$ where $N_{p}=NN_{\text{tor}}$ is the (integer) number of toroidal field periods within the off-set pattern and integer $M$ is the nearest to $\iota N_{p}$. Thus, the increasing sequence of $N_{p}$ corresponds to the sequence of the best Diophantine approximations $M/N_{p}$ of (generally irrational) $\iota$. These approximations are the convergents of $\iota$ representation by a simple continued fraction (see, e.g., Hardy and Wright (1985)). Therefore, an overall trend of the $N_{p}$ series $N_{p}^{(k)}$, $k=1,2,\dots$, is similar to that of geometrical progression, $N_{p}^{(k+1)}=b_{k}N_{p}^{(k)}$ with $b_{k}>1$ such that $b_{k}-1\gtrsim 1$. Thus, due to $\nu_{\ast}\propto N_{p}^{-5}$ following from Eq. (117), a series of collisionalities $\nu_{\ast}^{(k)}$ corresponding to the dominance of pattern $k$ in the off-set has similar geometrical trend, and a significant change in the off-set of $\lambda_{bB}$ requires such a change of $\log\nu_{\ast}$.
Finally, in view of the off-set (123) being of the same order as the collisionless asymptotic (42), let us estimate the role of the correction term driven by the source (67) which has been ignored in the present analysis and estimated to have a vanishing effect with decreasing collisionality in a short field line example in Section 3.4. Comparison of the asymptotic and numerical solutions in Fig. 25 shows an obvious feature which cannot be described by the leading order off-set. Namely, instead of the continuous linear growth of $g_{o}^{t}$, numerical solution tends to a nearly-periodic behaviour with the period of the order of the off-set pattern length. In other words, finite collisionality provides an aperiodic correction term which nearly balances the aperiodic behaviour of $g_{0}^{t}$ (while the leading order off-set provides only a nearly periodic correction contained in the off-set domains within the pattern). One can formally present this next order correction as a “staircase function”, $\Delta g_{\text{off}}^{\text{(tr)}}=\sum_{k}C_{k}^{\text{(tr)}}\Theta\left(\varphi-\varphi_{k}^{\text{(op)}}\right)$, where $\varphi_{k}^{\text{(op)}}$ are the boundaries of the off-set pattern (points of well aligned relevant maxima) and $C_{k}^{\text{(tr)}}=g_{0}^{t}\left(\varphi_{k-1}^{\text{(op)}}\right)-g_{0}^{t}\left(\varphi_{k}^{\text{(op)}}\right)$. Being of the same order as $g_{0}^{t}$ which, due to its dependence on $\varphi$, makes an independent of collisionality contribution to $\lambda_{bB}$ within each off-set pattern, the staircase function is constant within the pattern and, according to the first estimate (123), its contribution scales with collisionality. Namely, replacing the distribution function off-set $g_{\text{off}}^{(j)}$ with $\Delta g_{\text{off}}^{\text{(tr)}}\sim NN_{\text{pt}}$ where $N_{\text{pt}}$ is the number of patterns per closed field line (and $N$, as before, is the number of toroidal turns within the pattern), respective contribution to the bootstrap coefficient scales as $\Delta\lambda_{bB}\approx w_{\text{off}}^{(j)}\Delta g_{\text{off}}^{\text{(tr)}}\sim N_{\text{pt}}N^{-2}\sim N_{\text{pt}}\nu_{\ast}^{2/5}$.
Another feature demonstrated for the short field line in Fig. 8 is a split of the boundary layer caused by a strong alignment of relevant maxima. This split can also be formally described by a staircase function $\Delta g_{\text{off}}^{\text{(bl)}}=\sum_{k}C_{k}^{\text{(bl)}}\Theta\left(\varphi-\varphi_{k}^{\text{(op)}}\right)\delta_{\text{(b)}}(\eta_{b}-\eta)$ where $\delta_{\text{(b)}}$ tends with vanishing collisionality to a $\delta$-function and is localized within the boundary layer width $\delta\eta$ where it scales as $\delta\eta^{-1}$ at finite collisionality, whereas constants $C_{k}^{\text{(bl)}}$ are given by the integral in parentheses in Eq. (50) with the change of lower and upper limits to $\varphi_{k}^{\text{(op)}}$ and $\varphi_{k-1}^{\text{(op)}}$, respectively. Since $\Delta g_{\text{off}}^{\text{(bl)}}$ is also a constant of $\varphi$ within the pattern, its contribution to $\lambda_{bB}$ can be estimated as $\Delta\lambda_{bB}\sim\Delta w_{\text{off}}^{(j)}\Delta g_{\text{off}}^{\text{(bl)}}$ where $\Delta g_{\text{off}}^{\text{(bl)}}\sim NN_{\text{pt}}\delta\eta^{-1}$, and the estimate $\Delta w_{\text{off}}^{(j)}\sim N^{-1}\Delta I_{\cal V}^{(j)}$ follows from (111) where $\Delta I_{\cal V}^{(j)}\sim\delta\vartheta^{3}\sim N^{-3}$ is the contribution of the boundary layer to the integral $I_{\cal V}^{(j)}$ (this estimate is similar to the estimate of the error due to the finite $\eta_{\text{loc}}^{(j)}-\eta_{b}$ below Eq. (4.2)). Finally, we obtain $\Delta\lambda_{bB}\sim N_{\text{pt}}N^{-3}\delta\eta^{-1}\sim N_{\text{pt}}N^{-1}\sim N_{\text{pt}}\nu_{\ast}^{1/5}$ where we used (117) treated again as an approximate equality instead of a strong inequality. Thus, despite a visible effect on the generalized Spitzer function $g_{(3)}$, the ignored correction provides to the bootstrap coefficient a contribution which is converging to zero with decreasing collisionality.
It should be noted that scaling $\Delta\lambda_{bB}\propto N_{pt}$ with the number of off-set patterns within the closed field line is the result of the “worst case” estimate allowing for “bootstrap resonances” discussed below in Section 4.4. Such a crude “worst case” estimate of the collisionless asymptotic $\lambda_{bB}^{\dagger}$ in the form (54) results in a similar scaling with the field line length which, in turn, does not contain the collisionality dependence.